# MSGNav: Unleashing the Power of Multi-modal 3D Scene Graph for Zero-Shot Embodied Navigation

### v2

@@ Abstract
Embodied navigation is a fundamental capability for robotic agents operating. Real-world deployment requires open vocabulary generalization and low training overhead, motivating zero-shot methods rather than task-specific RL training. However, existing zero-shot methods that build explicit 3D scene graphs often compress rich visual observations into text-only relations, leading to high construction cost, irreversible loss of visual evidence, and constrained vocabularies. To address these limitations, we introduce the Multi-modal 3D Scene Graph (M3DSG), which preserves visual cues by replacing textual relational edges with dynamically assigned images. Built on M3DSG, we propose MSGNav, a zero-shot navigation system that includes a Key Subgraph Selection module for efficient reasoning, an Adaptive Vocabulary Update module for open vocabulary support, and a Closed-Loop Reasoning module for accurate exploration reasoning. Additionally, we further identify the “last mile” problem in zero-shot navigation — determining the feasible target location with a suitable final viewpoint, and propose a Visibility-based Viewpoint Decision module to explicitly resolve it. Comprehensive experimental results demonstrate that MSGNav achieves state-of-the-art performance on GOAT-Bench and HM3D-OVON datasets. The
open-source code will be publicly available at here.
@@ 1 Introduction
Embodied navigation is a fundamental capability of robotic agents, enabling them to accomplish diverse tasks in realistic environments [ovon, VLNsurvey]. Such tasks are typically goal-oriented, with targets defined by objects, natural language descriptions, or images [unigoal]. Recently, the challenging open-vocabulary embodied navigation task [mtu3d] has gained increasing attention, as real-world agents are required to generalize to a wide and unconstrained range of object categories [ovon].
Previous RL-based embodied navigation methods [rlnav1, rlnav2] suffer from poor generalization and a large sim-to-real gap [sg-nav]. Inspired by rapid advances in pre-trained large model techniques [gpt4o, qwen], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [3dmem, dynavlm, vlngame, tango] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
Notably, a series of methods [3dsgnav1, 3dsgnav2, saynav] that build explicit 3D scene graphs from observations and leverage LLMs to drive exploration have shown promising performance on standard goal-oriented navigation benchmarks [hm3d, mp3d]. However, traditional 3D scene graphs overly abstract 3D object relations into simple textual labels (e.g., “top”, “beside”), which seriously impedes progress on zero-shot embodied navigation. As illustrated in Fig. 1, the extreme abstracted 3D object-relation in traditional 3D scene graphs introduces three main limitations:
@@
Figure 2: Performance comparisons between our method and other existing ones for embodied navigation. (a) The superiority of our M3DSG over traditional 3D scene graphs. (b) Distance statistics from the goal for the previous method (3D-Mem [3dmem] as an example). (c) Our MSGNav system achieves state-of-the-art performance on both HM3D-OVON and Goat-Bench benchmarks.
@@
Additionally, we realize that the last‑mile problem in navigation tasks — knowing the target’s location is not equivalent to determining a suitable navigation pose [^1] Cannot see the wood for the trees. Being too close or occluded prevents a clear observation of the target, which is not a good viewpoint.. Existing zero-shot embodied navigation methods usually stop before that final step. As shown in Fig. 2(b), the existing 3D-Mem method [3dmem] successfully approaches the target but still fails in many navigation tasks due to unacceptable viewpoints (distances ranging from 0.25 m to 1.0 m). To address this issue, we introduce a visibility‑based viewpoint decision module in our MSGNav.
As shown in Fig. 2(c), MSGNav significantly outperforms state-of-the-art RL‑based and zero-shot methods on the Val‑Unseen set of challenging embodied navigation GOAT‑Bench [goatbench] and HM3D-OVON [ovon], demonstrating consistent and substantial gains in zero‑shot embodied navigation. Our contributions can be summarized as follows:
@@
  • MSGNav: We develop a zero-shot embodied navigation system built on M3DSG, enabling active exploration and achieving state-of-the-art performance on the GOAT-Bench and HM3D-OVON embodied navigation datasets.
@@ 2.1 Graph-based Scene Exploration
Graph representations [Conceptgraphs, related_14, related_33] provide explicit spatial–semantic structures for scene reasoning and integrate well with LLMs/VLMs. Examples include SayPlan [related_32] for 3D scene graph planning, OVSG [related_05] for open-vocabulary grounding, imaginative world modeling [hu2025imaginative], and BEV graph policies [luo2025learning]. Graph reasoning has also been applied in navigation, e.g. SG-Nav [sg-nav] for hierarchical prompting and graph-retained adaptation [hong2025general] for cross-domain VLN. Yet, the overly abstract textual edges of current scene graphs limit the comprehensive scene context understanding, motivating our multi-modal 3D scene graph (M3DSG).
@@ 2.2 Zero-Shot Navigation
Previous embodied navigation methods, primarily based on reinforcement learning (RL) [rlnav1, rlnav2], suffer from limited generalization and a significant sim-to-real gap, which hinders their application in real-world environments. Zero-shot navigation allows agents to act in unseen environments without task-specific training, unlike supervised methods [related_11, related_18, related_31, related_41] which require large-scale simulation. Existing zero-shot navigation methods span object-goal (ON), image-goal (IIN), and text-goal (TN). ON methods include CoW [related_12] with CLIP [related_28] frontier exploration, ESC [related_49], OpenFMNav [related_08], VLFM [vlfm] using LLM reasoning, UniGoal [unigoal] for multi-modal unification, and RATE-Nav [li2025rate] with region-aware termination. IIN methods range from Mod-IIN [imggoal2] to SIGN [yan2025sign] for safety-aware drone navigation. TN methods such as InstructNav [related_25] generalize instruction following across ON, DDN, and VLN [related_09, related_20, related_26, related_47, related_48]. However, most zero-shot work ignores the last-mile challenge of choosing an optimal viewpoint after target localization.
@@ 3 Approach
In this section, we formalize the zero-shot goal-oriented embodied navigation problem and provide an overview of our pipeline (Sec. 3.1). We then introduce the proposed Multi‑modal 3D Scene Graph (M3DSG) (Sec. 3.2), and describe MSGNav, the navigation system built on M3DSG (Sec. 3.3).
@@ 3.2 M3DSG: Multi‑modal 3D Scene Graph
Traditional 3D scene graphs represent object relations with text-only edges, which are overly abstract and hinder VLMs from accurately understanding scene contexts. Inspired by 3D-Mem [3dmem], which emphasizes the value of raw images for navigation, we retain visual information when constructing an object-centric 3D scene graph. Unlike traditional 3D scene graph [Conceptgraphs] which uses textual relation edges, our method stores images to describe detailed object relations directly. This image edge preserves the benefits of 3D scene graphs while avoiding repeated, costly model queries and providing a more holistic scene representation.
@@ 3.2.2 Incremental Construction of M3DSG
Following prior work [sg-nav, unigoal], the scene graph $\mathbf{S}$ is incrementally updated during exploration. At step $t$, $\mathbf{S}_{t-1}$ is updated to $\mathbf{S}_{t}$ with the current RGB-D observation $\mathcal{I}_{t}$ as:
@@
Object Update. For the RGB-D observation $\mathcal{I}_{t}$ at time $t$, we extract the frame-level object set $\mathbf{O}^{frame}_{t}=\{o_{i}\}_{i=1}^{|\mathbf{O}^{frame}_{t}|}$ by VFMs: YOLO-W [yolo] for open-vocabulary detection, SAM [sam] for instance masks, and CLIP [clip] for visual embeddings. Each object $o_{i}$ is represented as $o_{i}=\{\mathcal{ID},\mathcal{C},\mathcal{P},\mathcal{B},\mathcal{M},\mathcal{PC},\mathcal{V},\mathcal{R}\}_{i},$ denoting unique ID, category, 3D position, bounding box, semantic mask, point cloud, visual feature, and room location, respectively. The global object set $\mathbf{O}_{t}$ is incrementally updated from the previous set $\mathbf{O}_{t-1}$ and the frame-level object set $\mathbf{O}^{frame}_{t}$ as:
@@
where $\Phi_{\mathrm{match}}$ aligns objects in $\mathbf{O}^{frame}_{t}$ and $\mathbf{O}_{t-1}$ via spatial similarity $(\mathcal{B},\mathcal{PC})$ and visual similarity $(\mathcal{C},\mathcal{V})$, and $\Phi_{\mathrm{merge}}$ merges features of matched objects [Conceptgraphs]. This process maintains a global object set $\mathbf{O}_{t}$ within $\mathbf{S}_{t}$ step by step.
@@ 3.3.1 Key Subgraph Selection (KSS)
During exploration, the 3D scene graph $\mathbf{S}$ grows progressively and can become voluminous, thereby impairing the decision-making efficiency [3dmem]. In embodied navigation, however, typically only a small related subgraph to the required target is meaningful. Therefore, processing the full scene graph consumes substantial computational resources without commensurate benefits. The key to addressing this issue is to preserve maximal relevant information with the minimal scene graph. MGSNav extracts the target‑relevant subgraphs via a Compress–Focus–Prune procedure:
@@
where $\mathbf{F}$ denotes frontier images for exploration as in  [3dmem], $g$ is the navigation target, and $\mathcal{R}_{t}$ denotes the selected target or frontier for task response. The specific prompt design is provided in the supplementary material.
@@ 3.3.2 Adaptive Vocabulary Update (AVU)
Existing methods achieve open-vocabulary navigation simply by using detectors, e.g., YOLO-W [yolo] and GroundingDINO [groundingdino], but remain constrained by preset vocabulary (e.g., ScanNet-200 [scannet], HM3D [hm3d]). Such constraints hinder VLMs from handling diverse out-of-vocabulary scene representations in the real world. Additionally, preset vocabulary is difficult to accurately describe the image target.
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
Figure 4: Demonstration of the “last-mile” problem. (a) Previous methods select the nearest traversable position after target localization, and often fail due to poor or distant viewpoints. (b) Our VVD samples candidate viewpoints and computes visibility, and is able to select a suitable viewpoint close to the ground-truth (GT) viewpoint for successful navigation.
For a navigation agent, its high-level decision process involves the selection of a target coordinate, while the low-level planner converts it into executable actions. Therefore, an optimal final decision should choose a viewpoint with clear visibility to facilitate subsequent operations. Previous zero-shot methods [vlmnav, dynavlm, 3dmem, sg-nav] often assume that knowing the target location is enough, and select the nearest traversable position as the navigation viewpoint. However, as shown in Fig. 4(a), this may yield navigation viewpoints with occlusion or poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last-mile” problem of navigation—identifying a suitable viewpoint despite correctly locating the target $\bar{o}$. To solve this problem, we realize that a good viewpoint should have clear visibility of the target. To achieve this goal, we propose a Visibility-based Viewpoint Decision (VVD) module (in Algorithm 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation coordinate. Compared with traditional nearest-point choice, VVD consistently yields viewpoints with superior target visibility.
@@ 4.1 Experimental Setting
We evaluate on two goal-oriented open-vocabulary navigation benchmarks: HM3D-OVON [ovon] (open-vocabulary navigation), and GOAT-Bench [goatbench] (multi-modal lifelong open-vocabulary navigation). The metric of navigation task includes Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ indicates success of task, $l^{s}_{i}$ and $l^{a}_{i}$ represent shortest path length and agent path length, respectively. To demonstrate generalization capabilities, we evaluate approaches on the “Val Unseen” set of GOAT-Bench and HM3D-OVON datasets. We adopt GPT-4o (2024-08-06 version) as the choice of VLM by directly utilizing the OpenAI API. We also provide results using Qwen-VL-Max (2025-08-13 version) as the choice of VLM in the Appendix.

### v3

@@ Abstract
Embodied navigation is a fundamental capability for robotic agents operating. Real-world deployment requires open vocabulary generalization and low training overhead, motivating zero-shot methods rather than task-specific RL training. However, existing zero-shot methods that build explicit 3D scene graphs often compress rich visual observations into text-only relations, leading to high construction cost, irreversible loss of visual evidence, and constrained vocabularies. To address these limitations, we introduce the Multi-modal 3D Scene Graph (M3DSG), which preserves visual cues by replacing textual relational edges with dynamically assigned images. Built on M3DSG, we propose MSGNav, a zero-shot navigation system that includes a Key Subgraph Selection module for efficient reasoning, an Adaptive Vocabulary Update module for open vocabulary support, and a Closed-Loop Reasoning module for accurate exploration reasoning. Additionally, we further identify the “last mile” problem in zero-shot navigation — determining the feasible target location with a suitable final viewpoint, and propose a Visibility-based Viewpoint Decision module to explicitly resolve it. Comprehensive experimental results demonstrate that MSGNav achieves state-of-the-art performance on the challenging GOAT-Bench and HM3D-ObjNav benchmark. The
code will be publicly available at here.
@@ 1 Introduction
Embodied navigation is a fundamental capability of robotic agents, enabling them to accomplish diverse tasks in realistic environments [45, 52]. Such tasks are typically goal-oriented, with targets defined by objects, natural language descriptions, or images [43]. Recently, the challenging open-vocabulary embodied navigation task [55] has gained increasing attention, as real-world agents are required to generalize to a wide and unconstrained range of object categories [45].
Previous RL-based embodied navigation methods [28, 33] suffer from poor generalization and a large sim-to-real gap [42]. Inspired by rapid advances in pre-trained large model techniques [14, 1], zero-shot embodied navigation approaches have emerged at this historic moment. Zero-shot navigation system does not require any training or fine-tuning before being applied to real-world scenarios. These approaches [41, 15, 47, 56] substantially reduce training time and computational cost while enabling open‑vocabulary navigation, thereby facilitating deployment in real‑world applications.
Notably, a series of methods [36, 38, 30] that build explicit 3D scene graphs from observations and leverage LLMs to drive exploration have shown promising performance on standard goal-oriented navigation benchmarks [31, 2]. However, traditional 3D scene graphs overly abstract 3D object relations into simple textual labels (e.g., “top”, “beside”), which seriously impedes progress on zero-shot embodied navigation. As illustrated in Fig. 1, the extreme abstracted 3D object-relation in traditional 3D scene graphs introduces three main limitations:
@@
Figure 2: Performance comparisons between our MSGNav and other existing methods for embodied navigation on Goat-Bench [16]: the multi-modal open-vocabulary navigation benchmark. (a) The superiority of our M3DSG over traditional 3D scene graphs. (b) Distance statistics from the goal for the previous method (3D-Mem [41] as an example). (c) Our MSGNav system achieves state-of-the-art performance on Goat-Bench.
@@
Additionally, we realize that the last‑mile problem in navigation tasks — knowing the target’s location is not equivalent to determining a suitable navigation pose [^1] Cannot see the wood for the trees. Being too close or occluded prevents a clear observation of the target, which is not a good viewpoint.. Existing zero-shot embodied navigation methods usually stop before that final step. As shown in Fig. 2(b), the existing 3D-Mem method [41] successfully approaches the target but still fails in many navigation tasks due to unacceptable viewpoints (distances ranging from 0.25 m to 1.0 m). To address this issue, we introduce a visibility‑based viewpoint decision module in our MSGNav.
As shown in Fig. 2(c), MSGNav significantly outperforms state-of-the-art RL‑based and zero-shot methods on the challenging embodied navigation GOAT‑Bench [16], demonstrating consistent and substantial gains in zero‑shot challenging embodied navigation. Our contributions can be summarized as follows:
@@
  • MSGNav: We develop a zero-shot embodied navigation system built on M3DSG, enabling active exploration and achieving state-of-the-art performance on the GOAT-Bench and HM3D-ObjNav embodied navigation datasets.
@@ 2.1 Graph-based Scene Exploration
Graph representations [11, 17, 35] provide explicit spatial–semantic structures for scene reasoning and integrate well with LLMs/VLMs. Examples include SayPlan [34] for 3D scene graph planning, OVSG [3] for open-vocabulary grounding, imaginative world modeling [13], and BEV graph policies [26]. Graph reasoning has also been applied in navigation, e.g. SG-Nav [42] for hierarchical prompting and graph-retained adaptation [12] for cross-domain VLN. Yet, the overly abstract textual edges of current scene graphs limit the comprehensive scene context understanding, motivating our multi-modal 3D scene graph (M3DSG).
@@ 2.2 Zero-Shot Navigation
Previous embodied navigation methods, primarily based on reinforcement learning (RL) [28, 33], suffer from limited generalization and a significant sim-to-real gap, which hinders their application in real-world environments. Zero-shot navigation allows agents to act in unseen environments without task-specific training, unlike supervised methods [32, 39] which require large-scale simulation. Existing zero-shot navigation methods span object-goal (ON), image-goal (IIN), and text-goal (TN). ON methods include CoW [9] with CLIP [29] frontier exploration, ESC [54], OpenFMNav [20], VLFM [44] using LLM reasoning, UniGoal [43] for multi-modal unification, and RATE-Nav [22] with region-aware termination. IIN methods range from Mod-IIN [19] to SIGN [40] for safety-aware drone navigation. TN methods such as InstructNav [24] generalize instruction following across ON, DDN, and VLN [5, 21, 25, 48, 53]. However, most zero-shot work ignores the last-mile challenge of choosing an optimal viewpoint after target localization.
@@ 3 Approach
In this section, we formalize the zero-shot goal-oriented embodied navigation problem and provide an overview of our pipeline (Sec. 3.1). We then detail the construction of our Multi‑modal 3D Scene Graph (M3DSG) (Sec. 3.2), and describe MSGNav, the navigation system that performs effective and closed-loop reasoning over this graph (Sec. 3.3).
@@ 3.2 M3DSG: Multi‑modal 3D Scene Graph
Traditional 3D scene graphs represent object relations with text-only edges, which are overly abstract and hinder VLMs from accurately understanding scene contexts. Inspired by 3D-Mem [41], which emphasizes the value of raw images for navigation, we retain visual information when constructing an object-centric 3D scene graph. Unlike traditional 3D scene graph [11] which uses textual relation edges, our method stores images to describe detailed object relations directly. This image edge preserves the benefits of 3D scene graphs while avoiding repeated, costly model queries and providing a more holistic scene representation.
@@ 3.2.2 Incremental Construction of M3DSG
Following prior work [42, 43], the scene graph $\mathbf{S}$ is incrementally updated during exploration. At step $t$, $\mathbf{S}_{t-1}$ is updated to $\mathbf{S}_{t}$ with the current RGB-D observation $\mathcal{I}_{t}$ as:
@@
Object Update. For the RGB-D observation $\mathcal{I}_{t}$ at time $t$, we extract the frame-level object set $\mathbf{O}^{frame}_{t}=\{o_{i}\}_{i=1}^{|\mathbf{O}^{frame}_{t}|}$ by VFMs: YOLO-W [6] for open-vocabulary detection, SAM [18] for instance masks, and CLIP [29] for visual embeddings. Each object $o_{i}$ is represented as $o_{i}=\{\mathcal{ID},\mathcal{C},\mathcal{P},\mathcal{B},\mathcal{M},\mathcal{PC},\mathcal{V},\mathcal{R}\}_{i},$ denoting unique ID, category, 3D position, bounding box, semantic mask, point cloud, visual feature, and room location, respectively. The global object set $\mathbf{O}_{t}$ is incrementally updated from the previous set $\mathbf{O}_{t-1}$ and the frame-level object set $\mathbf{O}^{frame}_{t}$ as:
@@
where $\Phi_{\mathrm{match}}$ aligns objects in $\mathbf{O}^{frame}_{t}$ and $\mathbf{O}_{t-1}$ via spatial similarity $(\mathcal{B},\mathcal{PC})$ and visual similarity $(\mathcal{C},\mathcal{V})$, and $\Phi_{\mathrm{merge}}$ merges features of matched objects [11]. This process maintains a global object set $\mathbf{O}_{t}$ within $\mathbf{S}_{t}$ step by step.
@@ 3.3.1 Key Subgraph Selection (KSS)
During exploration, the 3D scene graph $\mathbf{S}$ grows progressively and can become voluminous, thereby impairing the decision-making efficiency [41]. In embodied navigation, however, typically only a small related subgraph to the required target is meaningful. Therefore, processing the full scene graph consumes substantial computational resources without commensurate benefits. The key to addressing this issue is to preserve maximal relevant information with the minimal scene graph. MGSNav extracts the target‑relevant subgraphs via a Compress–Focus–Prune procedure:
@@
where $\mathbf{F}$ denotes frontier images for exploration as in  [41], $g$ is the navigation target, and $\mathcal{R}_{t}$ denotes the selected target or frontier for task response. The specific prompt design is provided in the supplementary material.
@@ 3.3.2 Adaptive Vocabulary Update (AVU)
Existing methods achieve open-vocabulary navigation simply by using detectors, e.g., YOLO-W [6] and GroundingDINO [23], but remain constrained by preset vocabulary (e.g., ScanNet-200 [7], HM3D [31]). Such constraints hinder VLMs from handling diverse out-of-vocabulary scene representations in the real world. Additionally, preset vocabulary is difficult to accurately describe the image target.
@@ 3.3.4 Visibility-based Viewpoint Decision (VVD)
Figure 4: Demonstration of the “last-mile” problem. (a) Previous methods select the nearest traversable position after target localization, and often fail due to poor viewpoints. (b) Our VVD samples candidate viewpoints and computes visibility, which can select a suitable viewpoint close to GT for successful navigation.
For a navigation agent, the high-level decision process involves the location of the target, while the low-level planner converts it into executable actions. Therefore, an optimal final decision should choose a viewpoint with clear visibility to facilitate subsequent operations. Previous zero-shot methods [10, 15, 41, 42] often assume that knowing the target location is enough, and select the nearest traversable position as the navigation viewpoint. However, as shown in Fig. 4(a), this may yield viewpoints with poor visibility, causing failure even when the target $\bar{o}$ is correctly localized.
We term this the “last-mile” problem of navigation—identifying a suitable viewpoint despite correctly locating the target $\bar{o}$. To solve this problem, we realize that a good viewpoint should have clear visibility of the target. To achieve this goal, we propose a Visibility-based Viewpoint Decision (VVD) module (in Algorithm 2). As shown in Fig. 4(b), VVD uniformly samples candidate viewpoints around target $\bar{o}$ within radii $\mathbf{R}=\{r_{j}\}_{j=1}^{N_{R}}$. For each candidate $\mathbf{v}_{i}$, VVD computes a visibility score by evaluating occlusion between $\mathbf{v}_{i}$ and target point cloud $\mathcal{PC}_{\bar{o}}$. The viewpoint with the highest visibility score $\textbf{v}_{best}$ is selected as the navigation coordinate. Compared with traditional choices, VVD yields viewpoints with superior target visibility. Further explanation and analysis of the VVD module will be supplemented in the Appendix.
@@ 4.1 Experimental Setting
Datasets and Evaluation Metrics.
We evaluate our proposed approach on two established goal-oriented navigation benchmarks: GOAT-Bench [16], a multi-modal lifelong open-vocabulary dataset, and the widely adopted HM3D-ObjNav [31] (Habitat 2022 ObjectNav Challenge). To validate the generalization capability of our method to novel environments, all evaluations on GOAT-Bench are conducted on the Val Unseen split. Following standard practice, we assess navigation performance using Success Rate ($\text{SR}=\frac{N_{\text{success}}}{N_{\text{total}}}$) and Success weighted by Path Length ($\mathrm{SPL}=\frac{1}{N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}S_{i}\frac{l^{s}_{i}}{\max(l^{s}_{i},l^{a}_{i})}$), where $S_{i}\in\{0,1\}$ denotes binary task success, and $l^{s}_{i}$ and $l^{a}_{i}$ represent the shortest path distance and the actual path traversed by the agent.
Implementation Details.
We employ GPT-4o (version 2024-08-06) as our primary Vision-Language Model (VLM) via the OpenAI API. Additional analyses utilizing Qwen-VL-Max as an alternative VLM backbone, along with extended evaluations on the HM3D-OVON benchmark [45], are deferred to the supplementary material.